1. A stock price is currently $100. Over each of the next two six-month periods it is expected to go up by 10% or down by 10%. The risk-free rate is 8% per annum with continuous compounding.

(a) What is the value of a one-year European call option with a strike price of $100?

(b) What is the value of a one year European put option with a strike price of $100?

(c) What is the value of a one-year American call option with a strike price of $100?

(d) What is the value of a one-year American put option with a strike price of $100?
Answer #1

a

Step 5 Binomial Tree Pricing Step 1 Step 2 Scenario price=Spot price*(U)^2=100*1.1^2=121 Payoff HH=Max(Scenario price-Strike

b

Step 5 Binomial Tree Pricing Step 1 Step 2 Scenario price=Spot price*(U)^2=100*1.1^2=121 Payoff HH=Max(Strike price-Scenario

c

Binomial Tree Pricing Step 1 Step 5 Scenario pricesSpot price*(U)=100*1.1=110 Payoff H=Max(Scenario price-Strike price,0) =Ma

d

Binomial Tree Pricing Step 1 Step 5 Scenario price=Spot price*(U)=100*1.1=110 Payoff H=Max(Strike price-Scenario price, 0) =M